The question is complete, and please answer ALL of the boxes by the info provided. thanks

Short Straddle Short Straddle Composition: Short a call and a put with the same strike and expiration $35.00 $30.00 Max ProfiCalls Bide Ask 5.81€ 5.88 Puts Ask Bid 2. Straddle using options on TUV stock You wish to short the January 7.50 straddle. ThStrangles Strangles are very similar to straddles in many ways: they are composed of a combination of puts and calls, and forLong Strangle $4.00 Long Strangle on YZA (stock at $150) Long 1 YZA Oct 110 put @ $6.125 Long 1 YZA Oct 185 call @ $6.50 = LoPuts Calls Bid | Ask 139.00 139.06 Bid Ask Strangle using options on YZA stock You wish to go long one YZA Oct 90/210 stranglShort Strangle Short Strangle Composition: Short a put with a low strike price and short a call with a high strike price, samCalls Puts Strangle using options on EFG stock Bide Ask Bid 160.13 Aske 160.38 You wish to short three contracts for the Oct
Answer #1

2]

TUV Stock

1]

Net premium received for each straddle = Jan 7.50 call option premium + Jan 7.50 put option premium.

The premium received for selling any option is the bid price.

Net premium received for each straddle = $0.50 + $1.88 = $2.38

Net premium received for selling 3 straddles = $2.38 * 3 = $7.14

2]

Upper breakeven = strike price + net premium received for each straddle

Upper breakeven = $7.50 + $2.38 = $9.88

Lower breakeven = strike price - net premium received for each straddle

Lower breakeven = $7.50 - $2.38 = $5.12

3]

If the stock rallies to $11 :

The put option will expire worthless.

Loss on each call option = stock price at expiration - strike price = $11 - $7.50 = $3.50

Net premium received on each straddle = $2.38

Loss on each straddle = loss on each call option - net premium received on each straddle

Loss on each straddle = $3.50 - $2.38 = $1.12

Loss on 3 straddles = $1.12 * 3 = $3.36

4]

If the stock moves to $6 :

The call option will expire worthless.

Loss on each put option = strike price - stock price at expiration = $7.50 - $6 = $1.50

Net premium received on each straddle = $2.38

Profit on each straddle = net premium received on each straddle - loss on each put option

Profit on each straddle = $2.38 - $1.50 = $0.88

Loss on 3 straddles = $0.88 * 3 = $2.64

5]

If the stock falls to $4 :

The call option will expire worthless.

Loss on each put option = strike price - stock price at expiration = $7.50 - $4 = $3.50

Net premium received on each straddle = $2.38

Loss on each straddle = loss on each put option - net premium received on each straddle

Loss on each straddle = $3.50 - $2.38 = $1.12

Loss on 3 straddles = $1.12 * 3 = $3.36